OpenStudy (anonymous):

Can someone please explain this to me?? Let f(x) = -2x-7 and g(x) = -4x + 3. Find (f o g) (-5)

OpenStudy (anonymous):

the (x) plug in -5 and you should then be able to answer the problem if i am thinking correct

OpenStudy (owlcoffee):

When we deal a "composite function" notated "(f o g) (x)", we can read it like: "g(f(x))" So that means we "put function f inside function g". So, we know that: \[f(x)= -2x-7\] \[g(x)=-4x+3\] So, we want to find the composite function of f on g, so we'll look at function g, and when we see an "x", we will place "-2x-7" on it's place. Like this: \[g(f(x))=-4(-2x-7)+3\] Let's operate that: \[g(f(x))=8x+28+3\] \[g(f(x))=8x+31\] and now, that we have the composite function, we can evaluate on the point we were asked to evaluate it in, wich is x=-5. So let's take the function we just found and evaluate it: \[g(f(x))=8x+31\] then instead of "x" we'll put "-5": \[g(f(-5))=8(-5)+31\] doing some magical arithmetics: \[g(f(-5))=-40+31\] \[g(f(-5))=-9\] And voilá, that concludes the exercise. To resume what the steps are: 1) Find the "general" composite function, that means changing the "x" of one of the functions (careful with the order) and replace it with the function itself, confusing, I know, but you can always look at what I did above. 2) Evaluate the point in question, wich means, replacing the x's with the point.